OpenStudy (anonymous):

The first steps in writing f(x) = 4x^2 + 48x + 10 in vertex form are shown. f(x) = 4(x2 + 12x) + 10 What is the function written in vertex form?

OpenStudy (amorfide):

you want to complete the square \[x^{2}+bx\] to complete the square \[(x+\frac{ b }{ 2 })^{2} - (\frac{ b }{ 2 })^{2}\]

OpenStudy (amorfide):

then you would multiply your result by 4, then add on 10 to your constant then you have vertex form

OpenStudy (amorfide):

can you attempt this and tell me how you do? then I can help you if you get stuck

OpenStudy (anonymous):

how do i know what is x and b?

OpenStudy (amorfide):

x is just x the variable

OpenStudy (anonymous):

is 4 the value for b?

OpenStudy (amorfide):

b is the coefficient of x in your case \[x^{2} + 12x=x^{2} + bx\]

OpenStudy (amorfide):

so b=12

OpenStudy (amorfide):

any questions?

OpenStudy (anonymous):

i have no clue i keep getting mixed up

OpenStudy (amorfide):

okay so you took out a factor of 4 to start with so only focus on the inside of the brackets the inside of the brackets you have \[x^{2}+12x\]

OpenStudy (amorfide):

we will complete the square using the formula I gave you

OpenStudy (amorfide):

let b=12

OpenStudy (anonymous):

i tried that and i just got x

OpenStudy (anonymous):

no x squared

OpenStudy (amorfide):

so we get \[(x+\frac{ 12 }{ 2 })^{2} - (\frac{ 12 }{ 2 })^{2}\]

OpenStudy (amorfide):

this gives \[(x+6)^{2} - 6^{2}\]

OpenStudy (amorfide):

this gives \[(x+6)^{2} - 36\]

OpenStudy (anonymous):

4(x+6)^2 i don't know what to do with the rest

OpenStudy (amorfide):

okay now we have completed the square of what was inside the brackets we took out a factor of 4 earlier remember? so what we actually have is \[4\left[ (x+6)^{2}-36 \right] + 10\]

OpenStudy (amorfide):

multiply the bracket by 4 to get \[4(x+6)^{2} - 36 \times 4 + 10\]

OpenStudy (amorfide):

\[4(x+6)^{2} - 144 + 10\]

OpenStudy (amorfide):

simplify this :)

OpenStudy (amorfide):

is this okay? or do you have any questions?

OpenStudy (anonymous):

i think i got it. its -134

OpenStudy (amorfide):

yes so rewrite the whole thing to get \[4(x+6)^{2} - 134\] is your answer

OpenStudy (amorfide):

would appreciate a meddle for best answer if possible :)